OpenStudy (anonymous):

The midpoint M of KL has coordinates (12, 0.6). Point L has coordinates (14, 9). Find the coordinates of point K.

OpenStudy (anonymous):

tell me how to solve it and give me the answer

OpenStudy (jdoe0001):

\(\bf \textit{middle point of 2 points }\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ K&({\color{red}{ \square }}\quad ,&{\color{blue}{ \square }})\quad L&({\color{red}{ 14}}\quad ,&{\color{blue}{ 9}}) \end{array}\qquad \left(\cfrac{{\color{red}{ x_2}} + {\color{red}{ x_1}}}{2}\quad ,\quad \cfrac{{\color{blue}{ y_2}} + {\color{blue}{ y_1}}}{2} \right)\\ \quad \\\ \quad \\ \implies \left(\cfrac{{\color{red}{ 14}} + {\color{red}{ x_1}}}{2}\quad ,\quad \cfrac{{\color{blue}{ 9}} + {\color{blue}{ y_1}}}{2} \right)\quad =\quad (12,0.6)\implies \begin{array}{llll} \cfrac{{\color{red}{ 14}} + {\color{red}{ x_1}}}{2}=12\\ \quad \\ \bf \cfrac{{\color{blue}{ 9}} + {\color{blue}{ y_1}}}{2}=0.6 \end{array}\)